i AM LOOKING FOR THE THE EXCEL SET UP NOT INDIVIDUAL.

The production planner for Fine Coffees, Inc. produces two coffee blends: American (A) and British (B). He can only get 300 pounds of Colombian beans per week and 200 pounds of Dominican beans per week. Each pound of American blend coffee requires 12 ounces of Colombian beans and 4 ounces of Dominican beans, while a pound of British blend coffee uses 8 ounces of each type of bean. Profits for the American blend are $2.00 per pound, and profits for the British blend are $1.00 per pound. The goal of Fine Coffees, Inc. is to maximize profits.

Solve the LP problem in Excel and also perform sensitivity analysis.
Answer #1

EXCEL SETUP is following

FORMULAS:

American (A) British (B)
Result
Profit 2 1 =SUMPRODUCT(B5:C5,$B$3:$C$3)
Colombian beans 12 8 =SUMPRODUCT(B7:C7,$B$3:$C$3) <= =300*16 ounces 1 lb = 16 ounces
Dominican beans 4 8 =SUMPRODUCT(B8:C8,$B$3:$C$3) <= =200*16 ounces

Enter Solver Parameters

Click Solve button to generate the optimal solution

In the Solver results window, check Sensitivity and click OK to generate the sensitivity report

Solution:

Quantity of American (A) blend to be produced = 400 pounds

Quantity of British (B) blend to be produced = 0 pounds

Total profit = 2*400+1*0 = $ 800